to an 640870 , 11203333 and 1120*** (640870, 11203333) and (6407, 112015) are relatively prime. (040 TF
Answer #1

we have 640= 2^xs П1120 = 5x5x7 Рича .. - (uz) = (& sty) - , те 3333 16665 3333 3333 1 820 - 0 3333) 8 40 / Da o 6090 3 =from 0 and @ Q2) 10525 2463 3333 1 x 5 x 7 = (a,b)=1 =yes, a and b are relatively prime >Yes, given numbers are relatively pr